OpenStudy (thatonegirl_):

f(x)=x(4-x)^6 what are all of the critical points?

OpenStudy (tkhunny):

Have you considered the 1st Derivative?

OpenStudy (thatonegirl_):

Yeah that's where I'm having trouble. I used the product rule to find the derivative but idk how to find x lol

OpenStudy (tkhunny):

Demonstrate? What did the derivative look like?

OpenStudy (thatonegirl_):

(4-x)^6+x[-6(4-x)^5]

OpenStudy (tkhunny):

That's good. Now factor it and make it prettier. It will also be more convenient.

OpenStudy (thatonegirl_):

Factor or foil?

OpenStudy (tkhunny):

"FOIL" doesn't exist. Let's leave that in Algebra 1. Please FACTOR. Find common factors in each term and apply the Distributive Property.

OpenStudy (thatonegirl_):

okay so distribute

OpenStudy (thatonegirl_):

6 times..

OpenStudy (tkhunny):

No, factor. a*b + a*c = a*(b+c)

OpenStudy (thatonegirl_):

ahh so (4-x)^5(4-x-6x) ???

OpenStudy (tkhunny):

That's the spirit. Now what?

OpenStudy (thatonegirl_):

set them both equal to 0, but for the (4-x-6x), is that [(4-x)(-6x)] or (4-7x)?

OpenStudy (tkhunny):

You have to do better than that. 4 - x - 6x = ??? You tell me. No guessing.

OpenStudy (thatonegirl_):

well it's equal to 0. But when i factored out the (4-x) in the beginning, there was a -6x(4-x) left. So it should really be -24x+6x^2=0 right?

OpenStudy (tkhunny):

Please simplify properly, using addition, 4 - x - 6x = ??

OpenStudy (thatonegirl_):

why addition?

OpenStudy (tkhunny):

Do you see anything else? Combine like terms using whatever operation is there.

OpenStudy (thatonegirl_):

well i was saying before that it should be multiplied right?? not added together

OpenStudy (thatonegirl_):

OH WAIT NVM

OpenStudy (thatonegirl_):

I was looking at the wrong part of it. THANK YOU SO MUCH

OpenStudy (tkhunny):

Where is the multiplication in 4 - x - 6x? You once turned it into multiplication when you wrote "[(4-x)(-6x)]". However, that was only magic. You cannot turn addition (or subtraction) into multiplication.

OpenStudy (thatonegirl_):

So the anwers are x=4, 4/7

OpenStudy (thatonegirl_):

Yea i was factoring it out weird in my head :P

OpenStudy (tkhunny):

Not quite. Those are the x-values of the critical points. You have not quite answerd the "where are" part of the question.

OpenStudy (thatonegirl_):

oh well I would plug those back into the original f(x) to find the y. But this is multiple choice and it just says x= lol

OpenStudy (tkhunny):

Fair enough. Good work.

OpenStudy (thatonegirl_):

Thank you for the help and your time! :)